OpenStudy (anonymous):

I have 4 questions (CONICS SECTIONS! (HYPERBOLAS, PARABOLAS, ELLIPSES and CIRCLES!!!)) 1. For 9y^2-4x^2-18y+24x-63=0 What is the domain and range? 2. For y^2=8x, the vertex is (0,0, and focus is (2,0) and the directrix is x=-2 right? 3. (y-1)^2=-4(x+1), the vertex is (-1,1), the focus (0,1) and the directrix is x=2 right? 4. (x+2)^2=12y... what is the vertex, focus, directrix? Please help, thank you!

OpenStudy (anonymous):

Please help guys.. :( I attempted these and I just need someone to confirm it.

OpenStudy (tkhunny):

What did you get on #1 after you completed the square in both x and y?

OpenStudy (anonymous):

\[\frac{( y-1)^2 }{ 4 }-\frac{ (x-3)^2 }{ 9 }=1\]

OpenStudy (tkhunny):

Okay. I didn't check that. I'll just trust you. With the center of the hyperbola in hand, we'll need the vertices in order to determine the Domain and Range. What are the Vertices?

OpenStudy (anonymous):

Wait, the answer is (3, -3) and (3, 5)

OpenStudy (mertsj):

Number 3 opens to the left. Therefore the focus should be 1 unit to the left of the vertex.

OpenStudy (anonymous):

Is the vertex right?

OpenStudy (mertsj):

Yes. The focus is (-2,1) and the directrix is x = 0

OpenStudy (mertsj):

4. Vertex (-2,0) Focus (-2,3) directrix y = -3

OpenStudy (anonymous):

Thank you so much, you have no idea how much you saved me.